OpenStudy (anonymous):

cassie makes 2 candles candle 1 4cm wide, 6cm long, and 10cm high Candle2 8cm wide, 10cm long and 12cm high. write a ratio of the volume of candle 1 to candle 2

OpenStudy (anonymous):

Do you know how to find the volume of a cylindrical object?

OpenStudy (anonymous):

umm no

OpenStudy (anonymous):

You can assume the candles have rectangular shapes, or you can assume they have a cylindrical shape. Which one do you prefer?

OpenStudy (anonymous):

rectangle

OpenStudy (anonymous):

Volume for candle 1 is: V1=L*W*H = (6cm)(4cm)(10cm)=240 cubic centimeters. You try with Volume for candle 2

OpenStudy (anonymous):

8x10x12=960 for candle 2

OpenStudy (anonymous):

According with Wikipedia, a ratio is a relationship between two numbers of the same kind[1] (i.e., objects, persons, students, spoonfuls, units of whatever identical dimension), usually expressed as "a to b" or a:b, sometimes expressed arithmetically as a dimensionless quotient of the two[2] which explicitly indicates how many times the first number contains the second (not necessarily an integer).[3] So we can say that the ratio r =V1/V2 =240/960=1/4

OpenStudy (anonymous):

or you can say V1 to V2 is 1/4. Likewise, you can state the same relationship as: V1:V2=1/4. Both express the same.

OpenStudy (anonymous):

Is that what you want?

OpenStudy (anonymous):

yes thank you!

OpenStudy (anonymous):

be my fan :)